OpenStudy (anonymous):

2/3 + 3m/5 = 31/15 How do i do this problem?

OpenStudy (anonymous):

Step 1: Get common denominators.

OpenStudy (anonymous):

how do you do that im really sorry ive forgotten this stuff and the fact im home schooled doesnt help >.<

OpenStudy (texaschic101):

what is the common denominator for 3,5, and 15 ? What can 3,5, and 15 all go into evenly ?

OpenStudy (anonymous):

isnt it 30?

OpenStudy (texaschic101):

well, 30 is one of them, but it would be easier to find the lowest common denominator...there is one lower then 30.....can you find it ?

OpenStudy (anonymous):

give me a second to write this out

OpenStudy (anonymous):

15 :D

OpenStudy (texaschic101):

you can use 30 if you wish, but you will just have to reduce at the end. Yes...15 is the LCD. Now multiply the entire equation by 15. 15 (2/3) + 15(3m/5) = 15(31/15) now what can you do with that

OpenStudy (texaschic101):

you need help ?

OpenStudy (anonymous):

yes =.= i am so sorry i feel so stupid its been so long i havnt really praticed

OpenStudy (texaschic101):

once you multiply by 15, you should get : 10 + 9m = 31 (subtract 10 from both sides) 9m = 31 - 10 9m = 21 m = 21/9 or 2 3/9 which reduces to 2 1/3

OpenStudy (anonymous):

Thank you so so so much :D

OpenStudy (texaschic101):

no problem :)